Three applicants are to be choosen at random out of 4 boys & 6 girls.what is the probability of selecting 1)all girls 2)all boys 3)at least one boy

1
Expert's answer
2021-12-31T08:27:45-0500

There are "\binom {10}{3}=\frac{10!}{3!\cdot 7!}=\frac{10\cdot 9\cdot 8}{1\cdot 2\cdot 3}=120" ways of choosing 3 applicants.


1) There are "\binom{6}{3}=\frac{6!}{3!\cdot 3!}=\frac{4\cdot 5\cdot 6}{1\cdot 2\cdot 3}=20" ways of choosing 3 girls.

Then the probability of selecting all girls is "\frac{20}{120}=\frac{1}{6}" .


2) There are "\binom{4}{3}=\frac{4!}{3!\cdot 1!}=4" ways of choosing 3 boys.

Then the probability of selecting all boys is "\frac{4}{120}=\frac{1}{30}" .


3) Events “all girls” and “at least one boy” are opposite. Therefore, the probability of selecting at least one boy is "1-\frac{1}{6}=\frac{5}{6}" .


Answers: 1) "\frac{1}{6}" ; 2) "\frac{1}{30}" ; 3) "\frac{5}{6}" .